# Advancing GPU IPC for Stiff Affine-Deformable Simulation

### v2

@@ 4.2. Connectivity-Enhanced MAS Construction
6:     $[domain\_Idx,domain\_Idy]=\mathbf{Map}(r,c)$triangleright align METIS and MAS subdomain with padding (Figure 6)
7:     $m=domain\_Idy/N$triangleright get MAS subdomian id
8:     $l=0$triangleright MAS level, starting from 0
9:     if $domain\_Idx/N==domain\_Idy/N$ then $\triangleright$ if nodes $r$ and $c$ belong to the same level-0 MAS subdomain, write $\mathbf{M3}$ to the level-0 MAS matrix
10:         $\mathbf{M}_{0,MAS}(m)[domain\_Idx\%N,domain\_Idy\%N]=\mathbf{M3}$
11:     $\triangleright$ add $\mathbf{M3}$ to higher level MAS matrix if needed
12:         for $l=1,2,...,levels-1$ do
13:              if $r$ and $c$ in the same subdomain at level $l$ then
14:                  AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
15:              end if
16:         end for
17:     end if
18:end for
19:for $BId$ $=$ 00, $1$, …, $len(\mathbf{M}_{0,MAS})-1$ do in parallel $\triangleright$ accumulate all $3\times 3$ blocks in $\mathbf{M}_{0,MAS}$ by warp reduction and add it to all higher-level MAS matrices
20:     $\mathbf{M3}={warpReduction}(\mathbf{M}_{0,MAS}(BId))$
21:     for $l=1,2,...,levels-1$ do
22:         AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
23:     end for
@@
25:return $\mathbf{M}_{MAS}^{-1}$
4.3. GPU Warp Reduction for Connectivity-Enhanced MAS
@@
In our case, the nodes within the same level-0 subdomain map to a small number (often the same) super nodes in coarser levels. This allows for the application of a warp reduction method to accumulate the level-0 matrix blocks. The accumulated matrix is then used to construct the sub-matrix of the next coarser levels using atomic operations. This approach effectively reduces the cost of MAS matrix precomputation, since significantly less atomic operations are needed as the level-1 matrix can be efficiently computed using warp reduction. This also allows for more levels in the MAS preconditioner with lower computational cost. See Algorithm 1 and our supplemental document for more details.
5. Cubic Strain-Limiting energy with Analytic eigensystem
We develop our cubic strain-limiting energy based on the FEM Baraff-Witkin (FBW) constitutive model (Kim, 2020). Inspired by Li et al. (2021) which augments soft membrane energy with a barrier-based strain-limiting term to avoid membrane locking, we aim to develop a barrier-free strain-limiting energy with analytic eigensystems so that expensive backtracking-based line search filtering for the strain limit and numerical eigendecomposition are both avoided. We will first introduce FBW and our model (subsection 5.1), and then derive the analytic eigensystem of our Hessian matrix (subsection 5.2).
@@ 5.1. Formulation
We follow Kim (2020) and use $\sqrt{I_{5}(\mathbf{F},\mathbf{n})}$ to define our strain-limiting energy, as it effectively represents the stretch factor in the direction $\mathbf{n}$. Directly incorporating $I_{5}$ into the barrier-based strain-limiting model of Li et al. (2021) enables reusing the eigenanalysis from Kim (2020). However, backtracking-based line search filtering, which is not parallel-friendly, is still required since no analytic bound for feasible step sizes is available for the strain. Additionally, because the barrier function diverges rapidly near the strain limit, solver performance becomes highly sensitive to the barrier stiffness, complicating parameter setting. To address this, we apply a cubic penalty function to handle the inequality constraint imposed by the strain limit. This approach allows for slight constraint violations while providing stretching resistance on the same scale as practical settings.
To simplify the explanation, let’s consider the $\mathbf{n}_{u}$ direction in Equation 9 as an example. We construct a $C^{2}$-continuous clamped cubic strain-limiting energy using $I_{5}$:
@@
and so the strain-limiting force is only applied when the cloth is stretched. By similarly defining the cubic term in the $\mathbf{n}_{v}$ direction, we obtain our complete cubic strain-limiting energy
@@
Figure 8. Our composed strain energy in the $\mathbf{n}_{u}$ direction. Combined with the membrane energy from Kim (2020), our cubic strain-limiting energy enables using soft compression resistance and stiff stretching resistance to avoid membrane locking and over-elongation issues. Here, $\lambda a_{t}=1$ and $\lambda^{{}^{\prime}}=100\lambda$ are used for visualization.
This energy is well-defined even if the strain limit is exceeded, eliminating the need for line search filtering. Given that the Young’s modulus of real fabrics ranges from $1MPa$ to $10MPa$ (Penava et al., 2014) when the cloth is stretched, setting our $\lambda^{\prime}=5MPa$ provides stretching resistance at the same scale when the cloth approaches the strain limit, e.g., $5\%$. Thus, our model prevents over-elongation in practical cloth manipulation scenarios. For $\Psi_{stretch}$, we can now safely use a small stiffness ($\lambda=0.05MPa$ in all our examples) to avoid membrane locking. Refer to Figure 8 for a plot of our composed membrane energy in the $\mathbf{n}_{u}$ direction. Since the stiffness for the shearing term $\Psi_{shear}$ primarily depends on fabric type and, based on our observations, has minimal impact on the over-elongation or membrane locking issues, we simply set it to 30% of the stretching stiffness and apply the original eigenanalysis from Kim (2020) to compute its PSD Hessian.
@@ 5.2. Eigenanalysis
Following Kim et al. (2019); Kim (2020), we can analytically derive the eigensystem of our $I_{5}$-based strain-limiting energy’s Hessian matrix with respect to the deformation gradient $\mathbf{F}$. For simplicity, we will omit $\lambda^{\prime}$ and $a_{t}$ in the following discussion.
@@
From these equations, we see that all eigenvalues are positive when $I_{5}>1$, indicating that our cubic strain-limiting energy $\Psi_{SL}$ is convex with respect to $\mathbf{F}$ (recall that $\Psi_{SL}=0$ when $I_{5}\leq 1$). Since $\mathbf{F}$ is a linear function of the degrees of freedom $x$, $\Psi_{SL}$ is also convex with respect to $x$.
@@ 6. Fast global hessian operations on the GPU
In this section, we introduce our multi-layer reduction strategy for fast Hessian matrix assembly in affine-deformable coupling simulations (subsection 6.1). Using a custom hash-based parallel reduction algorithm (subsection 6.2), our strategy significantly reduces the number of numerical operations and write conflicts, and it enables the development of a memory-efficient symmetric sparse blockwise matrix-vector multiplication method (subsection 6.3) to further boost PCG performance.
6.1. Multi-Layer Reduction for ABD Hessian Assembly
@@
However, the number of mapped Hessians typically far exceeds the number of affine body pairs (see Figure 9), leading to significant accumulation overhead, let alone the frequent write conflicts. To address this issue, we first accumulate all $3\times 3$ contact Hessian blocks with the same node index pair $\{r,c\}$ through reduction, then apply a single product with $\{\mathbf{J}_{r},\mathbf{J}_{c}\}$ to transform it into a $12\times 12$ matrix. This process yields a sequence of $12\times 12$ matrices for the blocks of the global Hessian corresponding to the affine body pair $\{AF(r),AF(c)\}$. A second layer of reduction is then used to accumulate these $12\times 12$ Hessians in parallel, significantly reducing write conflicts and the number of values to be accumulated, resulting in faster performance.
@@ 6.2. Hash-based Parallel Reduction
The first two primitives are standard in parallel computing. For warp-level segmented reduction, we use the CUB intrinsic function cub::WarpReduce(values, tags). Consider a simple example with a warp size of 8: if $\emph{{values}}=[1,1,1,1,1,1,1,1]$ and $\emph{{tags}}=[1,0,0,1,0,0,1,0]$, the output will be $\emph{{outPut}}=[3,\bullet,\bullet,3,\bullet,\bullet,2,\bullet]$, where $\bullet$ indicates invalid entries. Here, the tags array specifies start and end positions for reduction in values, with 1 marking the start. Reduction results (valid entries in outPut) are then written to the result array. The overall procedure, FastHashReduction, is detailed in Algorithm 2. First, the map array is constructed in parallel, mapping duplicate keys to their unique key indices to ensure reduction results are written to the correct locations (lines 1–8). The FastSegmentReduction algorithm (Algorithm 3) then efficiently computes the matrix.
@@
7:$UK,NK\leftarrow$ RunLengthEncode($K$) $\triangleright$ get unique keys ($UK$) and their counts ($NK$)
8:$Num\leftarrow$ len(UK)
9:$ESum\leftarrow$ ExclusiveSum($NK$) $\triangleright$ get prefix sum, starting from 0
10:// assign the unique key index to $O$ for each key in $K$:
11:$P\leftarrow$ Zeros(len(K))
12:for $tid=0,1,...,Num-1$ do in parallel
13:     $P[ESum[tid]+NK[tid]-1]\leftarrow 1$
14:end for
15:$O\leftarrow$ ExclusiveSum($P$) $\triangleright$ map each $K$ to unique key
16:$R\leftarrow$ FastSegmentReduction($O$,$V$)
@@
In Algorithm 2, we create a mapping array $O$ to associate each key-value pair with its target memory location, as illustrated in Figure 10. Array $O$ is also used to generate tag information for the CUB intrinsic function cub::WarpReduce(values, tags). The tags are determined by comparing adjacent entries in $O$, as described in Algorithm 3. Here, $b^{-}$, $b$, and $b^{+}$ represent the previous, current, and next unique key indices, respectively. The IsValid flag checks whether the GlobalThreadId falls within valid data range. If $b^{-}\neq b$, it marks the start of a new segment. The IsCrossWarp flag indicates if the segment spans multiple warps, necessitating atomic operations to avoid data corruption; otherwise, direct writing is safe. These flags are set per thread, but only the leading thread (head thread) can write to the output array. Therefore, we merge flags across threads in each segment: a head thread is valid if any thread in the segment is valid and marked as cross-warp if any spans multiple warps. This is achieved via segmented warp reduction, using an OR operation for merging (see lines 5-28).
@@
8:define WD = WARP_DIM
9:BlockCount $\leftarrow$ (len(V) + BD - 1) / BD
10:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD-1  do in parallel
11:     // get tags for cub::WarpReduce(values, tags):
12:     LaneId $\leftarrow$ GlobalThreadId % WD
13:     $b^{-},b,b^{+}\leftarrow-1$triangleright $-1$ for invalid value
14:     IsValid, IsHead, IsCrossWarp $\leftarrow$ 0
15:     Value $\leftarrow$ $\mathbf{0}$
16:     if $0<$ GlobalThreadId $<$ len(V) then
17:         $b^{-}\leftarrow O[$GlobalThreadId 1]-1]- 1 ]
18:     end if
19:     if GlobalThreadId $<$ len(V) - 1 then
20:         $b^{+}\leftarrow O[$GlobalThreadId +1]+1]+ 1 ]
@@
22:     if GlobalThreadId $<$ len(V) then
23:         $b\leftarrow O[$GlobalThreadId]]]]
24:         Value $\leftarrow V[$GlobalThreadId]]]]
25:         IsValid $\leftarrow 1$
26:     end if
27:     if LaneId ==0==0= = 0 then
28:         IsHead $\leftarrow 1$, IsCrossWarp $\leftarrow b^{-}==b$
29:     else
30:         IsHead $\leftarrow b^{-}\neq b$
31:         if LaneId $=$ WARP_DIM $-1$ then
32:              IsCrossWarp $\leftarrow b^{+}==b$
33:         end if
34:     end if
35:     Tuple $\leftarrow$ (IsValid, IsCrossWarp)
36:     Tuple $\leftarrow$ cub::WarpReduce(Tuple, IsHead)
37:     Value $\leftarrow$ cub::WarpReduce(Value, IsHead)
38:     // write the result to the corresponding memory location:
39:     if IsHead and IsValid then
40:         if IsCrossWarp then
41:              AtomicAdd($R[b]$, Value)
42:         else
43:              $R[b]\leftarrow$ Value
44:         end if
45:     end if
46:end for
@@ 6.3. Symmetric Reduce-By-Key (SRBK) SpMV
Sparse matrix-vector multiply (SpMV) is a frequently called subroutine in scientific computing, where a conjugate gradient solver, for example, may invoke hundreds or thousands of SpMV operations for each linear solve. Thus, speeding up SpMV is critical for improving the overall performance. The NVIDIA cuSparse library’s SpMV on Block Sparse Row (BSR) format is the fastest for general sparse matrices, but it lacks support for symmetric matrices. Since the SpMV algorithm is memory-bound, we aim to leverage symmetry to reduce memory access and achieve significant speedup.
@@
8:define WD = WARP_DIM
9:BlockCount $\leftarrow$ (len($\mathbf{AU}$) + BD - 1) / BD
10:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD11-1- 1  do in parallel
11:     Run line 5-29 in 3, replacing $O$ with $\mathbf{Rid}$
12:     if IsValid then $\triangleright$ GlobalThreadId is within valid data range
13:         j $\leftarrow\mathbf{Cid}[$GlobalThreadId]]]]
14:         $\mathbf{H}\leftarrow\mathbf{AU}[$GlobalThreadId]]]]
15:         Value $\leftarrow\mathbf{H}\cdot\mathbf{V}_{input}[j]$triangleright multiply with upper-triangular and diagonal blocks
16:         if $b\neq j$ then $\triangleright$ If not diagonal
17:              AtomicAdd($\mathbf{V}_{output}[j]$, $\mathbf{H}^{T}\cdot\mathbf{V}_{input}[b]$) $\triangleright$ multiply with lower-triangular blocks
18:         end if
19:     end if
20:     Run line 28-37 in 3, replacing $R$ with $\mathbf{V}_{output}$
21:end for

### v3

@@ 4.3. GPU Warp Reduction for Connectivity-Enhanced MAS
6:     $[domain\_Idx]=\mathbf{Map}(r)$
7:     $[domain\_Idy]=\mathbf{Map}(c)$triangleright align METIS and MAS subdomain with padding (Figure 6)
8:     $m=domain\_Idy/N$triangleright get MAS subdomain id
9:     $l=0$triangleright MAS level, starting from 0
10:     if $domain\_Idx/N==domain\_Idy/N$ then $\triangleright$ if nodes $r$ and $c$ belong to the same level-0 MAS subdomain, write $\mathbf{M3}$ to the level-0 MAS matrix
11:         $\mathbf{M}_{0,MAS}(m)[domain\_Idx\%N,domain\_Idy\%N]=\mathbf{M3}$
12:     $\triangleright$ add $\mathbf{M3}$ to higher level MAS matrix if needed
13:         for $l=1,2,...,levels-1$ do
14:              if $r$ and $c$ in the same subdomain at level $l$ then
15:                  AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
16:              end if
17:         end for
18:     end if
19:end for
20:for $BId$ $=$ 00, $1$, …, $len(\mathbf{M}_{0,MAS})-1$ do in parallel $\triangleright$ accumulate all $3\times 3$ blocks in $\mathbf{M}_{0,MAS}$ by warp reduction and add it to all higher-level MAS matrices
21:     $\mathbf{M3}={warpReduction}(\mathbf{M}_{0,MAS}(BId))$triangleright the warpReduction here utilizes the method of Gao et al. (2018)
22:     for $l=1,2,...,levels-1$ do
23:         AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
@@
25:end for
26:return $\mathbf{M}_{MAS}^{-1}$triangleright invert unpadded parts
@@
In our case, the nodes within the same level-0 subdomain map to a small number (often the same) super nodes in coarser levels. This allows for the application of a warp reduction method to accumulate the level-0 matrix blocks. The accumulated matrix is then used to construct the sub-matrix of the next coarser levels using atomic operations. This approach effectively reduces the cost of MAS matrix precomputation, since significantly less atomic operations are needed as the level-1 matrix can be efficiently computed using warp reduction. This also allows for more levels in the MAS preconditioner with lower computational cost. See Algorithm 1 and our supplemental document for more details. For meshes with fewer than 16 nodes, our method may group nodes from different meshes to minimize the number of subdomains. In such special scenarios, we revert to the original MAS assembly method in Wu et al. (2022).
5. Cubic Inexact Strain-Limiting energy with Analytic eigensystem
We develop our cubic inexact strain-limiting energy based on the FEM Baraff-Witkin (FBW) constitutive model (Kim, 2020). Inspired by Li et al. (2021) which augments soft membrane energy with a barrier-based strain-limiting term to avoid membrane locking, we aim to develop a barrier-free inexact strain-limiting energy with analytic eigensystems so that expensive backtracking-based line search filtering for the strain limit and numerical eigendecomposition are both avoided. We will first introduce FBW and our model (subsection 5.1), and then derive the analytic eigensystem of our Hessian matrix (subsection 5.2).
@@ 5.1. Formulation
Figure 8. Our composed strain energy in the $\mathbf{n}_{u}$ direction. Combined with the membrane energy from Kim (2020), our cubic inexact strain-limiting energy enables using soft compression resistance and stiff stretching resistance to avoid membrane locking and over-elongation issues. Here, $\lambda a_{t}=1$ and $\lambda^{{}^{\prime}}=100\lambda$ are used for visualization.
We follow Kim (2020) and use $\sqrt{I_{5}(\mathbf{F},\mathbf{n})}$ to define our inexact strain-limiting energy, as it effectively represents the stretch factor in the direction $\mathbf{n}$. Directly incorporating $I_{5}$ into the barrier-based strain-limiting model of Li et al. (2021) enables reusing the eigenanalysis from Kim (2020). However, backtracking-based line search filtering is still required since no analytic bound for feasible step sizes is available for the strain. This method is inefficient for parallelization because its iterative halving process, applied to each triangle, requires the global linear search iterations to continue until all triangles satisfy the strain limit. Additionally, because the barrier function diverges rapidly near the strain limit, solver performance becomes highly sensitive to the barrier stiffness, complicating parameter setting. To address this, we apply a cubic penalty function to inexactly handle the inequality constraint imposed by the strain limit. This approach allows for slight constraint violations while providing stretching resistance on the same scale as practical settings.
To simplify the explanation, let’s consider the $\mathbf{n}_{u}$ direction in Equation 9 as an example. We construct a $C^{2}$-continuous clamped cubic inexact strain-limiting energy using $I_{5}$:
@@
and so the inexact strain-limiting force is only applied when the cloth is stretched. This idea aligns with the tension field theory (Steigmann, 1990; Skouras et al., 2014; Chen et al., 2023), which deactivates stiffness under compression to obtain a convex energy. Additionally, Kikuuwe et al. (2009) activates a cubic term under compression to mitigate the element inversion issue.
Equivalently, Equation 10 can also be written as
$\Psi_{SL}^{u}=\lambda^{\prime}a_{t}\max(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1,0)^{3}.$
By similarly defining the cubic term in the $\mathbf{n}_{v}$ direction, we obtain our complete cubic inexact strain-limiting energy
@@
This energy is well-defined even if the strain limit is exceeded, eliminating the need for line search filtering. Given that the Young’s modulus of real fabrics ranges from $1MPa$ to $10MPa$ (Penava et al., 2014) when the cloth is stretched, we set our $\lambda^{\prime}=5MPa$ by default, as it provides stretching resistance at the same scale when the cloth approaches the strain limit, e.g., $5\%$. Thus, our model prevents over-elongation in practical cloth manipulation scenarios. For $\Psi_{stretch}$, we can now safely use a small stiffness ($\lambda=0.05MPa$ in all our examples) to avoid membrane locking. Refer to Figure 8 for a plot of our composed membrane energy in the $\mathbf{n}_{u}$ direction. Since the stiffness for the shearing term $\Psi_{shear}$ primarily depends on fabric type and, based on our observations, has minimal impact on the over-elongation or membrane locking issues, we simply set it to 30% of the stretching stiffness and apply the original eigenanalysis from Kim (2020) to compute its PSD Hessian.
@@ 5.2. Eigenanalysis
Following Kim et al. (2019); Kim (2020), we can analytically derive the eigensystem of our $I_{5}$-based inexact strain-limiting energy’s Hessian matrix with respect to the deformation gradient $\mathbf{F}$. For simplicity, we will omit $\lambda^{\prime}$ and $a_{t}$ in the following discussion.
@@
From these equations, we see that all eigenvalues are positive when $I_{5}>1$, indicating that our cubic inexact strain-limiting energy $\Psi_{SL}$ is convex with respect to $\mathbf{F}$ (recall that $\Psi_{SL}=0$ when $I_{5}\leq 1$). Since $\mathbf{F}$ is a linear function of the degrees of freedom $x$, $\Psi_{SL}$ is also convex with respect to $x$.
@@ 6. Fast global hessian operations on the GPU
In this section, we introduce our two-level reduction strategy for fast Hessian matrix assembly in affine-deformable coupling simulations (subsection 6.1). Specifically, using a custom hash-based parallel reduction algorithm (subsection 6.2), our strategy significantly reduces the number of numerical operations and write conflicts, and it enables the development of a memory-efficient symmetric sparse blockwise matrix-vector multiplication method (subsection 6.3) to further boost PCG performance.
6.1. Two-Level Reduction for ABD Hessian Assembly
@@
However, the number of mapped Hessians typically far exceeds the number of affine body pairs, which becomes even more pronounced in the affine-deformable coupling cases (see Figure 9), leading to significant accumulation overhead, let alone the frequent write conflicts. To address this issue, we first accumulate all $3\times 3$ contact Hessian blocks with the same node index pair $\{r,c\}$ through reduction, then apply a single product with $\{\mathbf{J}_{r},\mathbf{J}_{c}\}$ to transform it into a $12\times 12$ matrix. This process yields a sequence of $12\times 12$ matrices for the blocks of the global Hessian corresponding to the affine body pair $\{AF(r),AF(c)\}$. A second layer of reduction is then used to accumulate these $12\times 12$ Hessians in parallel, significantly reducing write conflicts and the number of values to be accumulated, resulting in faster performance.
@@ 6.2. Hash-based Parallel Reduction
The first two primitives are standard in parallel computing. For warp-level segmented reduction, we use the CUB [^2] https://nvidia.github.io/cccl/cub/ intrinsic function HeadSegmentedReduce(values, tags). Consider a simple example with a warp size of 8: if $\emph{{values}}=[1,1,1,1,1,1,1,1]$ and $\emph{{tags}}=[1,0,0,1,0,0,1,0]$, the output will be $\emph{{outPut}}=[3,\bullet,\bullet,3,\bullet,\bullet,2,\bullet]$, where $\bullet$ indicates invalid entries. Here, the tags array specifies start and end positions for reduction in values, with 1 marking the start. Reduction results (valid entries in outPut) are then written to the result array. The overall procedure, FastHashReduction, is detailed in Algorithm 2. First, the map array is constructed in parallel, mapping duplicate keys to their unique key indices to ensure reduction results are written to the correct locations (lines 1–8). The FastSegmentReduction algorithm (Algorithm 3) then efficiently computes the matrix.
@@
7:$UK\leftarrow$ RunLengthEncode($K$) $\triangleright$ get unique keys ($UK$), which are used to recover the row/column indices in the global Hessian
8:$Num\leftarrow$ len(K)
9:// assign the unique key index to $O$ for each key in $K$:
10:for $tid=0,1,...,Num-2$ do in parallel
11:     $P[i]\leftarrow K[i]\neq K[i+1]$
12:end for
13:     $O\leftarrow$ ExclusiveSum($P$) $\triangleright$ map each $K$ to unique key
14:     $R\leftarrow$ FastSegmentReduction($O$,$V$)
@@
In Algorithm 2, we create a mapping array $O$ to associate each key-value pair with its target memory location, as illustrated in Figure 10. Array $O$ is also used to generate tag information for the CUB intrinsic function HeadSegmentedReduce(values, tags). The tags are determined by comparing adjacent entries in $O$, as described in Algorithm 3. Here, $b^{-}$, $b$, and $b^{+}$ represent the previous, current, and next unique key indices, respectively. The IsValid flag checks whether the GlobalThreadId falls within valid data range. If $b^{-}\neq b$, it marks the start of a new segment. The IsCrossWarp flag indicates if the segment spans multiple warps, necessitating atomic operations to avoid data corruption; otherwise, direct writing is safe. These flags are set per thread, but only the leading thread (head thread) can write to the output array. Therefore, we merge flags across threads in each segment: a head thread is valid if any thread in the segment is valid and marked as cross-warp if any spans multiple warps. This is achieved via segmented warp reduction, using an OR operation for merging (see lines 5-28).
@@
8:BlockCount $\leftarrow$ (len(V) + BD - 1) / BD
9:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD-1  do in parallel
10:     // get tags for cub::WarpReduce(values, tags):
11:     $b^{-},b\leftarrow-1$triangleright $-1$ for invalid value
12:     LaneId $\leftarrow$ GlobalThreadId % WARP_DIM
13:     IsHead $\leftarrow$ 0
14:     Value $\leftarrow$ $\mathbf{0}$
15:     if GlobalThreadId $<$ len(O) then
16:         $b\leftarrow O[$GlobalThreadId]]]]
17:         Value $\leftarrow V[$GlobalThreadId]]]]
18:         if GlobalThreadId $>$then
19:              $b^{-}\leftarrow O[$GlobalThreadId 1]-1]- 1 ]
20:         end if
@@
22:     if LaneId ==0==0= = 0 Or $b^{-}\neq b$ then
23:         IsHead $\leftarrow 1$
24:     end if
25:     Value $\leftarrow$ HeadSegmentedReduce(Value, IsHead)
26:     // write the result to the corresponding memory location:
27:     if IsHead and GlobalThreadId $<$ len(O) then
28:         AtomicAdd($R[b]$, Value)
29:     end if
30:end for
@@ 6.3. Symmetric Reduce-By-Key (SRBK) SpMV
SpMV is a frequently called subroutine in scientific computing, where a conjugate gradient solver, for example, may invoke hundreds or thousands of SpMV operations for each linear solve. Thus, speeding up SpMV is critical for improving the overall performance. The NVIDIA cuSparse library’s SpMV on Block Sparse Row (BSR) format is the fastest for general sparse matrices, but it lacks support for symmetric matrices. Since the SpMV algorithm is memory-bound, we aim to leverage symmetry to reduce memory access and achieve significant speedup.
@@
8:BlockCount $\leftarrow$ (len($\mathbf{AU}$) + BD - 1) / BD
9:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD11-1- 1  do in parallel
10:     $b^{-},b\leftarrow-1$
11:     LaneId $\leftarrow$ GlobalThreadId % WARP_DIM
12:     IsHead $\leftarrow$ 0
13:     Value $\leftarrow$ $\mathbf{0}$
14:     if GlobalThreadId $<$ len($\mathbf{AU}$then
15:         $b\leftarrow\mathbf{Rid}[$GlobalThreadId]]]]
16:         j $\leftarrow\mathbf{Cid}[$GlobalThreadId]]]]
17:         $\mathbf{H}\leftarrow\mathbf{AU}[$GlobalThreadId]]]]
18:         Value $\leftarrow\mathbf{H}\cdot\mathbf{V}_{input}[j]$triangleright multiply with upper-triangular and diagonal blocks
19:         if $b\neq j$ then $\triangleright$ If not diagonal
20:              AtomicAdd($\mathbf{V}_{output}[j]$, $\mathbf{H}^{T}\cdot\mathbf{V}_{input}[b]$) $\triangleright$ multiply with lower-triangular blocks
21:         end if
22:         if GlobalThreadId $>$then
23:              $b^{-}\leftarrow\mathbf{Rid}[$GlobalThreadId1]-1]- 1 ]
24:         end if
25:     end if
26:     if LaneId ==0==0= = 0 Or $b^{-}\neq b$ then
27:         IsHead $\leftarrow 1$
28:     end if
29:     Value $\leftarrow$ HeadSegmentedReduce(Value, IsHead)
30:     // write the result to the corresponding memory location:
31:     if IsHead and GlobalThreadId $<$ len($\mathbf{AU}$then
32:         AtomicAdd($\mathbf{V}_{output}[b]$, Value)
33:     end if
34:end for